### v2

# Charges, complex structures, and perturbations of instantons

Abstract.
Hermitian non-Kähler Einstein 4-manifolds have a quasi-locally conserved charge associated to spin-lowering via Killing spinors, and corresponding to a parameter of the moduli space. This charge is evaluated for all explicitly known examples. Infinitesimal Einstein deformations are shown to admit a closed 2-form that measures the perturbation to this charge. Provided certain boundary conditions hold, a curve of metrics in the moduli space passing through a Hermitian non-Kähler Einstein metric is shown to be conformally Kähler to second perturbative order. This result is applied to prove that infinitesimal Ricci-flat deformations of ALF Hermitian non-Kähler instantons are tangent to the moduli space of Hermitian instantons, and hence as a consequence that infinitesimal rigidity holds. The analog of this statement also holds for the case of compact Hermitian non-Kähler Einstein metrics.
1. Introduction
In Euclidean quantum gravity, the dominant contribution to the gravitational path integral is expected to come from metrics near the minima of the Euclidean action [23], that is from gravitational instantons and small deformations of them. In the one-loop expansion, deformations of a classical solution $(M,g_{ab})$ represent quantum corrections [25], and in the case of infinitesimal Einstein deformations, i.e. perturbations satisfying the linearized Einstein equations, they provide information about the moduli space $\mathscr{E}(M)$ of instantons, describing the structure of the tangent space of $\mathscr{E}(M)$ at $g_{ab}$ and allowing to analyze questions such as integrability, infinitesimal rigidity [4] and spectral stability (e.g. [18, 11]); cf. [9] for the compact case. This paper is concerned with uncovering some special properties of this class of deformations.
The study of the moduli space is further motivated by the recent discovery of the Chen-Teo instanton [16], a Ricci-flat, asymptotically flat (AF) solution that gives a counter-example to the Euclidean black hole uniqueness conjecture [22, 24] and contributes, in addition to flat space and to the Schwarzschild and Kerr instantons, to the possible semi-classical states of the gravitational field with AF boundary conditions [32]. In principle, further unknown gravitational instantons could exist, and the classification problem remains largely open, similarly to the compact case where all known examples have special holonomy [9]; see [3].
For the vacuum Einstein equations with or without cosmological constant, other boundary conditions of interest in quantum gravity, in addition to AF, include asymptotically locally flat (ALF), asymptotically locally Euclidean (ALE), and compact manifolds. All known instantons are (at least locally [^1] Our definition of instantons here allows a non-vanishing cosmological constant. This includes the 4-sphere $S^{4}$, which does not admit a global complex structure.) Hermitian [2]. Furthermore, they are conformally Kähler, which means [19] that there is a metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$ and a self-dual 2-form $\hat{\omega}_{ab}$ which is parallel under the Levi-Civita connection ^$\hat{\nabla}_{a}$ of $\hat{g}_{ab}$,
^$\hat{\nabla}_{a}\hat{\omega}_{bc}=0.$ (1.1)
In this paper, we shall see that for Hermitian instantons, if the part of the Weyl tensor that has the same orientation as the complex structure, say $\Psi_{ABCD}$ in 2-spinor notation [39, 40] (cf. section 2.1 for conventions), is non-vanishing, then one of the moduli corresponds to a special “charge” associated to the gravitational field. This is because the Kähler form is actually the spin-lowered Weyl spinor
$\hat{\omega}_{ab}=\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}},$ (1.2)
where $K^{AB}$ is a real solution to the 2-index twistor equation, see section 3.1 below. The gravitational charges referred to above are then the periods of the closed 2-form (1.2) over 2-cycles $S$ in $M$,
$Q[S]=\int_{S}\hat{\omega}.$ (1.3)
Since ${\rm d}\hat{\omega}=0$, (1.3) is invariant under changes of representative of the homology class $[S]$, so the charge is both geometrical (it depends on the curvature of $g_{ab}$) and topological (it depends on the homology groups of $M$).
The above result is the Riemannian analog of a Lorentzian conservation law first found by Jordan, Ehlers, and Sachs [29], and then reformulated in [1] in terms of spin-lowering [40, §6.4] for the Kerr black hole, where the charge (1.3) corresponds to the mass parameter and the conservation law is intuitively clear: 2-spheres surrounding the black hole are non-contractible, and the mass contained within any two of them is the same. However, for gravitational instantons, the evaluation of (1.3) and its interpretation are considerably more subtle, since the topology may be substantially different from the Kerr topology $\mathbb{R}^{2}\times S^{2}$, and it is not obvious how to find the 2-cycles (i.e. the 2-spheres in Kerr). As we shall see, examples of (1.3) include: the NUT-charge/mass of the ALF Taub-bolt instanton on $\mathbb{CP}^{2}\setminus\{\rm pt.\}$, the only free parameter of the ALE Eguchi-Hanson instanton on $T^{*}\mathbb{CP}^{1}$, the cosmological constant for the compact instantons $\mathbb{CP}^{2}$ and $\mathbb{CP}^{2}\#\overline{\mathbb{CP}}^{2}$, and, remarkably, the two parameters of the AF Chen-Teo instanton on $\mathbb{CP}^{2}\setminus S^{1}$. The fact that the two parameters of Chen-Teo are integrals of the Kähler form is to be contrasted with the Kerr instanton, where (1.3) gives only the mass. Intuitively, the underlying reason is that the Chen-Teo instanton is in fact a “double” object (cf. [17, §5]), and each of the two 2-cycles carries a gravitational charge (1.3).
We shall prove that infinitesimal Einstein deformations of Hermitian instantons, with or without cosmological constant $\lambda$ and with any asymptotic structure, admit a 2-form $\delta\hat{\omega}_{ab}$ (given by (3.21)) which is closed, ${\rm d}\,\delta\hat{\omega}=0$, and whose integral over 2-cycles measures the perturbation to (1.3),
$\delta{Q}[S]=\int_{S}\delta\hat{\omega}.$ (1.4)
Moreover, the result (1.4) is gauge invariant (i.e. invariant under infinitesimal diffeomorphisms [^2] Gauge-invariance is important since $\mathscr{E}(M)$ involves a quotient by ${\rm Diff}(M)$.; see eq. (3.13) below). While (1.4) is valid for any boundary conditions, a particular application concerns the ALF Ricci-flat but non-Kähler case (the only possibilities being Kerr, Taub-bolt, and Chen-Teo [10, 35]), where infinitesimal Einstein deformations are expected to be perturbations with respect to the moduli parameters [4]. The formula (1.4) then gives the perturbation to one of the moduli of the Kerr, Taub-bolt, and Chen-Teo instantons.
The above result for perturbations is inspired by the conservation law for the Lorentzian case found by Fackerell [20] (cf. also [21]), which was formulated in terms of spin-lowering in [1] and used to argue that the analog of (1.4) gives the linearized mass of a perturbed Kerr black hole [1, Theorem 5.1].
Finally, while the conservation of (1.4) is the Riemannian analog of the Lorentzian result mentioned before, in this paper we shall also prove the following stronger result. For the compact and ALF cases, if $g_{ab}(s)$ is a smooth curve of Einstein metrics in $\mathscr{E}(M)$, with $g_{ab}(0)$ Hermitian, we shall prove that
$\left.\frac{{\rm d}}{{\rm d}{s}}\left[\hat{\nabla}_{a}^{(s)}\hat{\omega}_{bc}(s)\right]\right|_{s=0}=0.$ (1.5)
Here, we have a metric $\hat{g}_{ab}(s)=\Omega{(s)}^{2}g_{ab}(s)$ with Levi-Civita connection ^$\hat{\nabla}^{(s)}_{a}$, and $\hat{\omega}_{ab}(s)$ is a normalized eigenform of its self-dual Weyl tensor with eigenvalue $\Omega(s)$. In view of (1.1), the result (1.5) is gauge invariant. Note that, while the conservation law for (1.4) follows from (1.5) (by antisymmetrizing the indices), the result (1.5) is stronger.
The interpretation of (1.5) is as follows. The complex structure corresponding to $\hat{\omega}_{ab}(s)$ is parallel to second order, and hence the curve of metrics $g_{ab}(s)$ in $\mathscr{E}(M)$ is conformally Kähler and in particular Hermitian to second order, with charges $(Q+s\,\delta Q)[S]$. This is consistent with the recent result of Biquard, Gauduchon, and LeBrun [12], showing that local rigidity holds for Hermitian, toric, ALF instantons, i.e. that Ricci-flat metrics in a suitable neighbourhood of a Hermitian, toric, ALF Ricci-flat instanton are Hermitian. Local rigidity, however, does not imply infinitesimal rigidity, in the sense that an ALF Ricci-flat perturbation is tangent to the moduli space of Hermitian instantons. As an application of (1.5), we are now able to prove that infinitesimal rigidity holds.
Recall that an Einstein metric is said to be integrable if any linearized Einstein perturbation is tangent to a curve of Einstein metrics, see [9]. It is an open problem whether integrability holds in general for Einstein metrics. It was conjectured [4, Conjecture 1.7] that ALF Hermitian non-Kähler instantons are integrable in the above sense. From (1.5), we have that any infinitesimal Einstein deformation is Hermitian, and since the moduli space is smooth and finite-dimensional, we have the following theorem.
Theorem 1.1.
Let $(M,g_{ab})$ be an ALF Hermitian, non-Kähler gravitational instanton. Then $(M,g_{ab})$ is integrable and infinitesimal rigidity holds.
Overview
In section 2 we introduce some background material and set our notation and conventions. In section 3 we prove an identity for generic Riemannian 4-manifolds (Prop. 3.1), and use it to derive the closed 2-form $\delta\hat{\omega}_{ab}$ and the linearized charge (1.4); see Theorem 3.3. Section 4 contains many examples of the charge (1.3) for different gravitational instantons (AF, ALF, ALE, compact). In section 5, we prove (1.5) using a divergence identity (Lemma 5.1) based on LeBrun’s work [34], see Theorem 5.3, and in section 5.3 we use this to prove Theorem 1.1.
2. Preliminaries
2.1. Notation and conventions
Throughout the article, $(M,g_{ab})$ denotes a 4-dimensional, orientable Riemannian manifold, with metric signature (++++)(++++)( + + + + ), Levi-Civita connection $\nabla_{a}$, volume form $\epsilon_{abcd}$, and Hodge star operator *. We shall use, locally, the 2-spinor formalism [39, 40], adapted to Riemannian signature. We do not assume a global spin structure.
The (locally defined) unprimed and primed spinor bundles are $\mathbb{S}$ and $\mathbb{S}^{\prime}$. The symplectic forms $\epsilon_{AB}$ and $\epsilon_{A^{\prime}B^{\prime}}$ and their inverses are used to raise and lower spinor indices. An unprimed spin dyad is $(o^{A},\iota^{A})$ with $\epsilon_{AB}o^{A}\iota^{B}=1$ and $\iota^{A}=o^{\dagger A}$, and a primed spin dyad is $(\alpha^{A^{\prime}},\beta^{A^{\prime}})$ with $\epsilon_{A^{\prime}B^{\prime}}\alpha^{A^{\prime}}\beta^{B^{\prime}}=1$ and $\beta^{A^{\prime}}=\alpha^{\dagger A^{\prime}}$. Defining the null tetrad $\ell^{a}=o^{A}\alpha^{A^{\prime}}$, $n^{a}=\iota^{A}\beta^{A^{\prime}}$, $m^{a}=o^{A}\beta^{A^{\prime}}$, $\tilde{m}^{a}=\iota^{A}\alpha^{A^{\prime}}$ (see e.g. [39, (3.1.14)]), we have $n^{a}=\bar{\ell}^{a}$ and $\tilde{m}^{a}=-\bar{m}^{a}$, and the non-trivial inner products are $\ell^{a}n_{a}=1=-m^{a}\tilde{m}_{a}$. The real and imaginary parts of the null tetrad define four real vectors $\{e^{a}_{\bf a}\}$, ${\bf a}=0,...,3$, as follows: $e^{a}_{0}=\frac{1}{\sqrt{2}}(\ell^{a}+n^{a})$, $e^{a}_{1}=i\frac{1}{\sqrt{2}}(\ell^{a}-n^{a})$, $e^{a}_{2}=\frac{1}{\sqrt{2}}(m^{a}-\tilde{m}^{a})$, $e^{a}_{3}=i\frac{1}{\sqrt{2}}(m^{a}+\tilde{m}^{a})$. (This is the analog of [39, Eq. (3.1.20)].) The dual basis, denoted $e_{a}^{\bf a}$, is defined by $e_{a}^{\bf a}e^{a}_{\bf b}=\delta^{\bf a}_{\bf b}$. A computation gives:
$\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}=2(\ell_{(a}n_{b)}-m_{(a}\tilde{m}_{b)})=+\delta_{\bf ab}e_{a}^{\bf a}e_{b}^{\bf b}=+g_{ab}.$
Thus the reality conditions on the null tetrad imposed above, which are in turn defined by the Riemannian spinor conjugation $\dagger$, give the correct signature (++++)(++++)( + + + + ) for the metric. (Note that an analogous calculation but with Lorentzian reality conditions gives signature (+)(+---)( + - - - ), cf. footnote in [39, p. 121].)
Let $\Lambda^{2}$ be the space of 2-forms in $M$. This splits under * as $\Lambda^{2}=\Lambda^{2}_{+}\oplus\Lambda^{2}_{-}$, where $\Lambda^{2}_{\pm}$ corresponds to the eigenvalue ±$\pm 1$ of *. Elements of $\Lambda^{2}_{+}$ are self-dual 2-forms, and elements of $\Lambda^{2}_{-}$ are anti-self-dual 2-forms. Our convention for the volume form is $\epsilon_{0123}=+1$, i.e. $\epsilon_{abcd}=4!e^{0}_{[a}e^{1}_{b}e^{2}_{c}e^{3}_{d]}$, or equivalently $\boldsymbol{\epsilon}=+e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}$. With this convention, one can check that a basis of real self-dual 2-forms is:
$Z^{1}_{ab}=\tfrac{1}{\sqrt{2}}(e^{0}\wedge e^{3}+e^{1}\wedge e^{2})_{ab}=\tfrac{i}{\sqrt{2}}(o_{A}o_{B}-\iota_{A}\iota_{B})\epsilon_{A^{\prime}B^{\prime}},$ (2.1a)
$Z^{2}_{ab}=\tfrac{1}{\sqrt{2}}(e^{0}\wedge e^{2}-e^{1}\wedge e^{2})_{ab}=\tfrac{1}{\sqrt{2}}(o_{A}o_{B}+\iota_{A}\iota_{B})\epsilon_{A^{\prime}B^{\prime}},$ (2.1b)
$Z^{3}_{ab}=\tfrac{1}{\sqrt{2}}(e^{0}\wedge e^{1}+e^{2}\wedge e^{3})_{ab}=-\tfrac{i}{\sqrt{2}}(o_{A}\iota_{B}+\iota_{A}o_{B})\epsilon_{A^{\prime}B^{\prime}}.$ (2.1c)
Self-dual 2-forms are, thus, of the form $\omega^{+}_{ab}=\psi_{AB}\epsilon_{A^{\prime}B^{\prime}}$, while anti-self-dual 2-forms are $\omega^{-}_{ab}=\varphi_{AB}\epsilon_{A^{\prime}B^{\prime}}$. In other words, $\Lambda^{2}_{+}\cong\mathbb{S}\odot\mathbb{S}$ and $\Lambda^{2}_{-}\cong\mathbb{S}^{\prime}\odot\mathbb{S}^{\prime}$. Note that this convention is the opposite of [39, 40].
The 2-forms (2.1) give an orthonormal basis of $\Lambda^{2}_{+}$. In terms of spinors, this can be phrased in terms of the real ‘metric’ on $\mathbb{S}^{2}=\mathbb{S}\odot\mathbb{S}$ given by
$I_{ABCD}=\epsilon_{A(C}\epsilon_{|B|D)},$ (2.2)
cf. [40, Eq. (8.3.2)], with inverse $I^{ABCD}$. Writing (2.1) as $Z^{i}_{ab}=\zeta^{i}_{AB}\epsilon_{A^{\prime}B^{\prime}}$ ($i=1,2,3$), the spinors $\zeta^{i}_{AB}$ are real, symmetric, and satisfy
$I^{ABCD}\zeta^{i}_{AB}\zeta^{j}_{CD}=\delta^{ij},\qquad i=1,2,3.$ (2.3)
It also follows that $\delta_{ij}\zeta^{i}_{AB}\zeta^{j}_{CD}=I_{ABCD}$. Thus $\{\zeta^{i}_{AB}\}$ gives a real, orthonormal basis of $\mathbb{S}^{2}$.
We also note that the self-dual 2-forms $\omega^{i}_{ab}=\sqrt{2}Z^{i}_{ab}$ give (locally) an almost-hyper-Hermitian structure, i.e. they satisfy the quaternion algebra. In particular, each of $\omega^{i}_{ab}$ defines an almost-complex structure $J^{i}$ compatible with the metric, via $\omega^{i}=g(J^{i}\cdot,\cdot)$.
The Levi-Civita connection of $g_{ab}$ gives a connection in $\mathbb{S}^{2}$, also denoted $\nabla_{a}$. Since $\nabla_{a}:\Gamma(\mathbb{S}^{2})\to\Gamma(T^{*}M\otimes\mathbb{S}^{2})$, the covariant derivative of $\zeta^{i}_{BC}$ is
$\nabla_{a}\zeta^{i}_{BC}=\Gamma_{a}{}^{i}{}_{j}\zeta^{j}_{BC},$ (2.4)
where the connection 1-form on the right side is $\Gamma_{a}{}^{i}{}_{j}=\zeta_{j}^{BC}\nabla_{a}\zeta^{i}_{BC}$, and $\zeta_{j}^{BC}=\zeta^{jBC}$. The non-trivial, independent components are
$A_{a}={}$ $\Gamma_{a}{}^{3}{}_{1}=-(o^{B}\nabla_{a}o_{B}+\iota^{B}\nabla_{a}\iota_{B}),$ (2.5a)
$B_{a}={}$ $\Gamma_{a}{}^{3}{}_{2}=i(o^{B}\nabla_{a}o_{B}-\iota^{B}\nabla_{a}\iota_{B}),$ (2.5b)
$C_{a}={}$ $\Gamma_{a}{}^{2}{}_{1}=-2i\,\iota^{B}\nabla_{a}o_{B}.$ (2.5c)
Our convention for the Riemann tensor is
$(\nabla_{a}\nabla_{b}-\nabla_{b}\nabla_{a})v^{d}=-R_{abc}{}^{d}v^{c}.$ (2.6)
We stress that this is the opposite of the convention used by Penrose and Rindler [39, 40] (compare to [39, Eq. (4.2.31)]). This means that the Riemann tensor here is minus the Riemann tensor in [39, 40], and similarly the curvature spinors here are minus the curvature spinors in [39, 40]. This convention has been adopted to match the sign convention of the determinant of the Weyl tensor in references [42, 34, 12], which we shall invoke in section 5.
Using the convention (2.6), we have for any spinors $\kappa_{C},\chi_{C^{\prime}}$:
$\Box_{AB}\kappa_{C}={}$ $+\Psi_{ABC}{}^{D}\kappa_{D}+\frac{R}{24}(\epsilon_{AC}\kappa_{B}+\epsilon_{BC}\kappa_{A}),$ (2.7a)
$\Box_{A^{\prime}B^{\prime}}\chi_{C^{\prime}}={}$ $+\tilde{\Psi}_{A^{\prime}B^{\prime}C^{\prime}}{}^{D^{\prime}}\chi_{D^{\prime}}+\frac{R}{24}(\epsilon_{A^{\prime}C^{\prime}}\chi_{B^{\prime}}+\epsilon_{B^{\prime}C^{\prime}}\chi_{A^{\prime}}),$ (2.7b)
where $\Box_{AB}=\nabla_{A^{\prime}(A}\nabla^{A^{\prime}}_{B)}$ and $\Box_{A^{\prime}B^{\prime}}=\nabla_{A(A^{\prime}}\nabla^{A}_{B^{\prime})}$ (compare to [39, Eq. (4.9.15)]). In (2.7), $R=g^{ac}R_{abc}{}^{b}$ is the scalar curvature, and $\Psi_{ABCD}$ and $\tilde{\Psi}_{A^{\prime}B^{\prime}C^{\prime}D^{\prime}}$ are the Weyl curvature spinors. The Weyl tensor is $W_{abcd}=W^{+}_{abcd}+W^{-}_{abcd}$, where
$W^{+}_{abcd}=\Psi_{ABCD}\epsilon_{A^{\prime}B^{\prime}}\epsilon_{C^{\prime}D^{\prime}},\qquad W^{-}_{abcd}=\tilde{\Psi}_{A^{\prime}B^{\prime}C^{\prime}D^{\prime}}\epsilon_{AB}\epsilon_{CD}$ (2.8)
are, respectively, the self-dual and anti-self-dual parts (according to our conventions; cf. below (2.1)).
Finally, in section 3.1 we shall use a conformally invariant GHP connection $\mathcal{C}_{a}$ [6]. The necessary background is given in appendix A.
2.2. The self-dual Weyl spinor
In section 5 we shall be interested in eigenspinors for $\Psi_{ABCD}$. A detailed study of this can be found in [40, §8.3]. Here we collect some facts adapted to Riemannian signature, for later use in section 5.
Consider the map $\mathbb{S}^{2}\to\mathbb{S}^{2}$ given by
$\phi_{AB}\mapsto\Psi_{AB}{}^{CD}\phi_{CD}.$
(In tensor terms, $Z_{ab}\mapsto\frac{1}{2}W^{+}{}_{ab}{}^{cd}Z_{cd}$.) Choosing a real orthonormal basis $\{\zeta^{i}_{AB}\}$ of $\mathbb{S}^{2}$, the matrix $\Psi^{i}{}_{j}=\Psi_{AB}{}^{CD}\zeta^{i}_{CD}\zeta_{j}^{AB}$ is real and symmetric, so it can be diagonalized, and its eigenspaces are orthogonal w.r.t. $I_{ABCD}$. Denoting the corresponding real eigenspinors by $\phi^{i}_{AB}$, with (real) eigenvalues $\lambda_{i}$, we have
$\Psi_{AB}{}^{CD}\phi^{i}_{CD}=\lambda_{i}\phi^{i}_{AB}\qquad\text{(no sum over $i$)},$ (2.9)
where the $\phi^{i}_{AB}$’s are orthonormal, $I^{ABCD}\phi^{i}_{AB}\phi^{j}_{CD}=\delta^{ij}$. From [40, Eq. (8.3.8), (8.3.11)], the eigenvalues satisfy the following identities:
$\lambda_{1}+\lambda_{2}+\lambda_{3}={}$ $0,$ (2.10a)
$\lambda_{1}^{2}+\lambda_{2}^{2}+\lambda_{3}^{2}={}$ $\Psi_{AB}{}^{CD}\Psi_{CD}{}^{AB},$ (2.10b)
$3\lambda_{1}\lambda_{2}\lambda_{3}={}$ $\det(\Psi^{i}{}_{j}).$ (2.10c)
We see that $\Psi_{AB}{}^{CD}$ has at least one positive and one negative eigenvalue. We also note the following identities:
$\Psi_{ABCD}\phi_{i}^{AB}\phi_{j}^{CD}={}$ $\lambda_{i}\delta_{ij},$ (2.11a)
$\Psi_{ABCD}\nabla^{e}\phi^{AB}_{i}\nabla_{e}\phi_{j}^{CD}={}$ $\sum_{k=1}^{3}\lambda_{k}\Gamma^{ai}{}_{k}\Gamma_{a}{}^{j}{}_{k},$ (2.11b)
where $\Gamma_{a}{}^{i}{}_{j}$ is the connection 1-form defined as in (2.4) but w.r.t. the orthonormal basis of eigenspinors $\{\phi^{i}_{AB}\}$.
2.3. Conformal transformations
By a conformal transformation we mean a rescaling of the metric
$g_{ab}\to\hat{g}_{ab}=\Omega^{2}g_{ab},$ (2.12)
where $\Omega$ is a positive scalar field. The metric on self-dual 2-forms (2.2) transforms as $\hat{I}_{ABCD}=\Omega^{2}I_{ABCD}$. Since the Weyl spinor is conformally invariant, $\hat{\Psi}_{ABCD}=\Psi_{ABCD}$, then $\hat{\Psi}_{AB}{}^{CD}=\Omega^{-2}\Psi_{AB}{}^{CD}$. From (2.9), the eigenvalues of $\Psi_{AB}{}^{CD}$ transform as $\hat{\lambda}_{i}=\Omega^{-2}\lambda_{i}$. If $\phi^{i}_{AB}$ is an orthonormal basis of eigenspinors of $\Psi_{AB}{}^{CD}$, then a basis of eigenspinors of $\hat{\Psi}_{AB}{}^{CD}$ which is orthonormal w.r.t. $\hat{I}_{ABCD}$ is $\hat{\phi}^{i}_{AB}=\Omega\phi^{i}_{AB}$. Formulas (2.9), (2.10), (2.11) remain valid for the conformally transformed variables, after putting hats over all quantities. Consider one of the real self-dual 2-forms in (2.1), say $Z_{ab}\equiv Z^{3}_{ab}$. Then $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Under a conformal change (2.12), $\hat{Z}_{ab}=\Omega^{2}Z_{ab}$, so $J^{a}{}_{b}$ is conformally invariant. The Lee form $f=f_{a}{\rm d}{x}^{a}$ (see e.g. [28]) is defined by
${\rm d}{Z}=-2f\wedge Z.$ (2.13)
(In indices, $\partial_{[a}Z_{bc]}=-2f_{[a}Z_{bc]}$.) If $(M,g_{ab})$ is Einstein [^3] All of our results for Einstein manifolds apply in particular in the Ricci-flat case. We shall not mention this explicitly in the following. and Hermitian w.r.t. $J^{a}{}_{b}$, with $W^{+}_{abcd}\neq 0$, then the Goldberg-Sachs theorem implies that $Z_{ab}$ is an eigenform of $W^{+}{}_{ab}{}^{cd}$:
$\tfrac{1}{2}W^{+}{}_{ab}{}^{cd}Z_{cd}=-2\Psi_{2}Z_{ab},$ (2.14)
where $\Psi_{2}=-\frac{1}{8}W^{+}_{abcd}Z^{ab}Z^{cd}$. In this case, Bianchi identities imply that
$f_{a}=\Psi_{2}^{-1/3}\nabla_{a}\Psi_{2}^{1/3}.$ (2.15)
Since $\hat{g}_{ab}$ is also Hermitian, and the Lee form is (2.15), it follows that $\hat{g}_{ab}=\Psi_{2}^{2/3}g_{ab}$ is Kähler, and, defining $\omega_{ab}=\sqrt{2}Z_{ab}$, the (closed) Kähler form is
$\hat{\omega}_{ab}=\Psi_{2}^{2/3}\omega_{ab}.$ (2.16)
3. Linearized charges
3.1. Charges for Hermitian instantons
We shall extensively use the self-dual 2-form $Z^{3}_{ab}$ and spinor field $\zeta^{3}_{AB}$ in (2.1c), so it is convenient to denote
$\zeta_{AB}\equiv\zeta^{3}_{AB},\qquad Z_{ab}\equiv Z^{3}_{ab}.$ (3.1)
Proposition 3.1.
Let $(M,g_{ab})$ be a Riemannian manifold. Let $Z_{ab}=\zeta_{AB}\epsilon_{A^{\prime}B^{\prime}}$ be a real self-dual 2-form, with $\zeta_{AB}=-i\sqrt{2}o_{(A}\iota_{B)}$, and let $f_{AA^{\prime}}$ be the associated Lee form as in (2.13). Define
$\psi_{AB}=\Psi_{AB}{}^{CD}\zeta_{CD}.$ (3.2)
Then the following identity holds:
$\nabla_{A^{\prime}}{}^{B}\psi_{AB}=f_{A^{\prime}}{}^{B}\psi_{AB}+S_{AA^{\prime}}+K_{AA^{\prime}},$ (3.3)
where
$S_{AA^{\prime}}={}$ $\zeta^{CD}\nabla_{A^{\prime}}{}^{B}\Psi_{ABCD},$ (3.4)
$K_{AA^{\prime}}={}$ $i\sqrt{2}\,\Psi_{ABCD}(\sigma^{0}_{A^{\prime}}\iota^{B}\iota^{C}\iota^{D}+\sigma^{1}_{A^{\prime}}o^{B}o^{C}o^{D}),$ (3.5)
and $\sigma^{0}_{A^{\prime}},\sigma^{1}_{A^{\prime}}$ are defined in (A.7).
Remark 3.2.
The tensor version of (3.3) is
${\rm d}\Sigma=f\wedge\Sigma+*S+*K.$ (3.6)
where we defined the self-dual 2-form
$\Sigma_{ab}:=\Psi_{AB}{}^{CD}\zeta_{CD}\epsilon_{A^{\prime}B^{\prime}}=\tfrac{1}{2}W^{+}{}_{ab}{}^{cd}Z_{cd}.$ (3.7)
Proof.
The proof of (3.3) is straightforward if one uses a ‘conformal-GHP connection’ [6]: this is an operator $\mathcal{C}_{AA^{\prime}}$ which generalizes the ordinary GHP connection [39, §4.12] to be invariant under combined conformal transformations and rescalings of the spinors $o_{A},\iota_{A}$, by taking into account both the conformal and the GHP weights of the objects it acts on. The necessary background is given in appendix A. Since $\zeta_{AB}$ has conformal weight $w=1$, then $\zeta^{AB}$ has $w=-1$, and since $\Psi_{ABCD}$ has $w=0$, then $\psi_{AB}$ has $w=-1$. Thus:
$\nabla_{A^{\prime}}{}^{B}\psi_{AB}={}$ $\mathcal{C}_{A^{\prime}}{}^{B}\psi_{AB}$
[…]
where in the first line we used (A.5), in the third we used (A.6), and in the fourth we used (A.7) and the definition (3.5). ∎
If $(M,g_{ab})$ is Einstein, then the Weyl spinor satisfies
$\nabla^{AA^{\prime}}\Psi_{ABCD}=0,$ (3.8)
so the term with $S$ in (3.6) vanishes. Suppose that, in addition, $(M,g_{ab})$ is Hermitian, with complex structure $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ (so that the principal spinors are $o_{A},\iota_{A}$, recall (3.1) and (2.1c)). Then from appendix A, we have $\sigma^{0}_{A^{\prime}}=0$ and the term with $*K$ in (3.6) vanishes. Recalling that for an Einstein-Hermitian geometry with $\Psi_{2}\neq 0$, the Lee form is (2.15), the result (3.6) reduces to
${\rm d}\left[\Psi_{2}^{-1/3}\Sigma\right]=0.$ (3.9)
One way of interpreting this identity is in terms of spin-lowering via Killing spinors [40, §6.4]. We have $\Psi_{2}^{-1/3}\Sigma_{ab}\propto\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$, where $K_{AB}=\Psi_{2}^{-1/3}o_{(A}\iota_{B)}$ is a valence-2 Killing spinor, $\nabla_{A^{\prime}(A}K_{BC)}=0$ (cf. [40, Eq. (6.7.15)]). Then (3.9) is the statement that
$\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$ (3.10)
is an anti-self-dual Maxwell field. As explained by Penrose and Rindler in [40, §6.4], this construction can be applied to linearized gravity on Minkowski space-time by replacing $\Psi_{ABCD}$ with a spin 2 zero-rest-mass field. Then one can associate a charge to this field by integrating (3.10) on a closed 2-surface, where, since Minkowski space-time is topologically trivial, the 2-surface must surround a region of matter sources in order for the charge to be non-vanishing.
For our Einstein Hermitian geometry $(M,g_{ab})$, an analogous non-vanishing charge requires the existence of non-contractible 2-surfaces in $M$. Additionally, we can also interpret (3.9) in terms of complex structures: using (3.7) and (2.14), eq. (3.9) is a reflection of the fact that $(M,g_{ab})$ is conformally Kähler, with the Kähler form given by (2.16). Combining the two interpretations above, we can then define a charge associated to every 2-cycle $S$ in $M$, by formula (1.3). Note that this charge has a sort of combination of geometrical and topological nature, since it depends on both the curvature of $g_{ab}$ and the existence of non-trivial 2-cycles in $M$.
3.2. Perturbations and closed 2-forms
Consider a smooth one-parameter family of metrics $g_{ab}(s)$ in $M$. A tensor field $T(s)$ constructed from the metric has an expansion
$T(s)=T+(\delta{T})\,s+\frac{1}{2!}(\delta^{2}{T})\,s^{2}+O(s^{3}),$ (3.11)
where $T,\delta{T},\delta^{2}{T},...$ are tensor fields in $M$,
$\delta^{k}T=\left.\left[\frac{{\rm d}^{k}}{{\rm d}s^{k}}T(s)\right]\right|_{s=0}.$ (3.12)
We say that $T=T(s)\big{|}_{s=0}$ is the “background” value of $T(s)$, and $\delta{T}$ is the linearization of $T(s)$ at $T$. We recall that, under an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{a}$, the linearization $\delta{T}$ transforms as
$\delta{T}\to\delta{T}+\pounds_{\xi}T,$ (3.13)
which in this context is referred to as a gauge transformation. We say that the linearization $\delta{T}$ is gauge invariant if it is invariant under (3.13) for all $\xi^{a}$. Note that, in particular, this holds if the background value of $T(s)$ vanishes, i.e. $T=0$.
The Levi-Civita connection of $g_{ab}(s)$ is denoted $\nabla^{(s)}_{a}$, and its relation to the background Levi-Civita connection is (see [41, Eq. (7.5.7)])
$\nabla^{(s)}_{a}v_{b}=\nabla_{a}v_{b}-C_{ab}{}^{c}(s)\,v_{c}$ (3.14)
for all $v_{a}$, where
$C_{ab}{}^{c}(s)=\frac{1}{2}g^{cd}(s)\left[\nabla_{a}g_{bd}(s)+\nabla_{b}g_{ad}(s)-\nabla_{d}g_{ab}(s)\right]$ (3.15)
and $\nabla_{a}=\nabla_{a}^{(0)}$. Note that $C_{ab}{}^{c}(s)=s\,\delta{C}_{ab}{}^{c}+O(s^{2})$.
Theorem 3.3.
Let $g_{ab}(s)$ be a one-parameter family of Riemannian metrics in $M$, such that, for all $s$,
$\nabla^{(s)\,a}W^{+}_{abcd}(s)=0.$ (3.16)
Let $Z_{ab}(s)$ be a real self-dual 2-form, $Z_{ab}(s)Z^{ab}(s)=2$, and define
$\Sigma_{ab}(s):={}$ $\frac{1}{2}W^{+}{}_{ab}{}^{cd}(s)Z_{cd}(s),$ (3.17)
$\Psi_{2}(s):={}$ $-\frac{1}{8}W^{+}_{abcd}(s)Z^{ab}(s)Z^{cd}(s),$ (3.18)
$\hat{\omega}_{ab}(s):={}$ $-\frac{\sqrt{2}}{3}\left[\Psi_{2}(s)^{-1/3}\Sigma_{ab}(s)-\Psi_{2}(s)^{2/3}Z_{ab}(s)\right].$ (3.19)
Suppose the background metric $g_{ab}=g_{ab}(0)$ is Einstein and Hermitian, with (closed) self-dual Kähler form $\hat{\omega}_{ab}=\hat{\omega}_{ab}(0)$, and $\Psi_{2}=\Psi_{2}(0)\neq 0$. Then:
${\rm d}\,\delta\hat{\omega}=0,$ (3.20)
where $\delta\hat{\omega}$ is the 2-form
$\delta\hat{\omega}_{ab}=-\frac{\sqrt{2}}{3}\left[\Psi_{2}^{-1/3}\delta\Sigma_{ab}-\Psi_{2}^{2/3}\delta{Z}_{ab}\right].$ (3.21)
Moreover, the result (3.20) is gauge invariant.
Proof.
For all $s$, we have
${\rm d}{Z}(s)={}$ $-2f(s)\wedge Z(s),$ (3.22)
${\rm d}\Sigma(s)={}$ $f(s)\wedge\Sigma(s)+(*K)(s),$ (3.23)
where the first equality is simply the definition of the Lee form (recall (2.13)), and the second equality follows from (3.6) after using that (3.16) holds for all $s$. Since the background is Einstein and Hermitian, then $(*K)(s)=O(s^{2})$, so ${\rm d}(*K)/{\rm d}s\,|_{s=0}=0$. Thus, the linearization of (3.22)-(3.23) is
${\rm d}\,\delta Z={}$ $-2\delta{f}\wedge Z-2f\wedge\delta Z,$
${\rm d}\,\delta\Sigma={}$ $\delta{f}\wedge\Sigma+f\wedge\delta\Sigma,$
where $Z=Z(0)$, $\Sigma=\Sigma(0)$ and $f=f(0)$. Using the background identities (2.15) and (2.14), it follows that
$\delta{f}\wedge\Sigma=-2\Psi_{2}\delta{f}\wedge Z=\Psi_{2}({\rm d}\delta Z+2f\wedge\delta Z)=\Psi_{2}^{1/3}{\rm d}\left[\Psi_{2}^{2/3}\delta Z\right],$
so
${\rm d}\delta\Sigma=\Psi_{2}^{1/3}{\rm d}\left[\Psi_{2}^{2/3}\delta Z\right]+\Psi_{2}^{-1/3}{\rm d}\Psi_{2}^{1/3}\wedge\delta\Sigma$
and therefore
$\Psi_{2}^{1/3}{\rm d}\left[\Psi_{2}^{-1/3}\delta\Sigma\right]=\Psi_{2}^{1/3}{\rm d}\left[\Psi_{2}^{2/3}\delta Z\right],$
which gives (3.20).
Finally, consider a gauge transformation $\delta\hat{\omega}\to\delta\hat{\omega}+\pounds_{\xi}\hat{\omega}$. We have £$\pounds_{\xi}\hat{\omega}={\rm d}(\xi\lrcorner\,\hat{\omega})$ since $\hat{\omega}$ is closed. Hence, $\delta\hat{\omega}\to\delta\hat{\omega}+{\rm d}(\xi\lrcorner\,\hat{\omega})$, so the result (3.20) is gauge-invariant. ∎
4. Examples
In this section we shall see several examples of the charges (1.3), by computing the conformal Kähler structures, the 2-cycles, and then the periods (1.3), for both compact and non-compact instantons.
A non-trivial issue is how to find the independent 2-cycles, for which we need the second homology group $H_{2}(M)$. To do this, we shall use the fact that all of the examples are toric. This implies that one can apply the rod structure formalism as in [15], which allows to determine the topological properties of the instanton. For example, for a rod structure with $n$ turning points, the Euler characteristic $\chi(M)$ is equal to $n$. In the non-compact case, we see from [36, Theorem 1] that $H_{2}(M)\cong\mathbb{Z}^{n-1}$. In the compact case, we can use the fact that $\chi(M)$ is also the alternating sum of the Betti numbers $b_{i}(M)$ to deduce the number of independent 2-cycles.
Below we shall also use the terminology of ‘nuts’ and ‘bolts’, for which we refer to the paper by Gibbons and Hawking [24].
4.1. Ambi-Kähler examples
The examples in this subsection are conformally Kähler for both orientations. In [5], these geometries are referred to as ambi-Kähler. The complex structures are $\omega^{+}_{ab}=2io_{(A}\iota_{B)}\epsilon_{A^{\prime}B^{\prime}}$ and $\omega^{-}_{ab}=2i\tilde{o}_{(A^{\prime}}\tilde{\iota}_{B^{\prime})}\epsilon_{AB}$, and the Kähler forms are $\hat{\omega}^{\pm}_{ab}$. We denote the only non-trivial components of $W^{\pm}_{abcd}$ by $\Psi_{2}^{\pm}$. (Note $\Psi_{2}^{+}=\Psi_{2}$.)
Example 4.1 (Kerr).
The AF Kerr instanton with parameters $m>0,a$ has the following local form of the metric:
$g={}$ $\frac{\Delta}{\Sigma}({\rm d}\tau-a\sin^{2}\theta{\rm d}\phi)^{2}+\frac{\sin^{2}\theta}{\Sigma}[a{\rm d}\tau+(r^{2}-a^{2}){\rm d}\phi]^{2}+\frac{\Sigma}{\Delta}{\rm d}{r}^{2}+\Sigma{\rm d}\theta^{2},$ (4.1)
where $\Delta=r^{2}-2mr-a^{2}$, $\Sigma=r^{2}-a^{2}\cos^{2}\theta$. Both $W^{\pm}_{abcd}$ are type D, with
$\Psi_{2}^{\pm}=\frac{m}{(r\pm a\cos\theta)^{3}}.$ (4.2)
The following 2-forms
$\omega^{\pm}=({\rm d}\tau-a\sin^{2}\theta{\rm d}\phi)\wedge{\rm d}{r}\pm\sin\theta{\rm d}\theta\wedge[a{\rm d}\tau+(r^{2}-a^{2}){\rm d}\phi]$ (4.3)
define integrable almost-complex structures with opposite orientation. The Kähler forms are then given by $\hat{\omega}^{\pm}=(\Psi_{2}^{\pm})^{2/3}\omega^{\pm}$. The topology is $\mathbb{R}^{2}\times S^{2}$, and the rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.4]. Since there are two turning points, we have $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1], so there is only one independent 2-cycle $S$, corresponding to the bolt at $r=r_{+}$ of the Killing field $\partial_{\tau}+\Omega_{E}\partial_{\phi}$, where $r_{+}=m+\sqrt{m^{2}+a^{2}}$ and $\Omega_{E}=a/(2mr_{+})$. The associated charges (1.3) are
[…]
For our next two examples, we first recall that the Ricci-flat Taub-NUT metric, with parameters $m>0,n>0$ is
$g={}$ $4n^{2}\frac{\Delta}{\Sigma}({\rm d}\psi+\cos\theta{\rm d}\phi)^{2}+\frac{\Sigma}{\Delta}{\rm d}{r}^{2}+\Sigma({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}),$ (4.5)
where $\Sigma=r^{2}-n^{2}$, $\Delta=r^{2}-2mr+n^{2}$. Both $W^{\pm}_{abcd}$ are generically type D,
$\Psi_{2}^{\pm}=\frac{m\pm n}{(r\pm n)^{3}}.$ (4.6)
The 2-forms
$\omega^{\pm}=2n({\rm d}\psi+\cos\theta{\rm d}\phi)\wedge{\rm d}{r}\pm\Sigma\sin\theta{\rm d}\theta\wedge{\rm d}\phi$ (4.7)
correspond to integrable almost-complex structures $J^{\pm}$ with opposite orientation ($(J_{\pm})^{a}{}_{b}=\omega^{\pm}_{bc}g^{ca}$). These structures are not Kähler.
Example 4.2 (Self-dual Taub-NUT).
The self-dual Taub-NUT instanton is obtained from (4.5) when $m=n$. It is ALF, and the topology is $S^{4}\setminus\{\rm pt.\}\cong\mathbb{R}^{4}$. From (4.6), we see that $\Psi^{-}_{2}=0$ and $\Psi^{+}_{2}=2n/(r+n)^{3}$, that is $W^{-}_{abcd}=0$ and $W^{+}_{abcd}\neq 0$, hence the space is hyper-Kähler with respect to one orientation, i.e. there is a 2-sphere of Kähler structures, and type D with respect to the other. The 2-forms (4.7) are however not Kähler; the geometry is not only hyper-Kähler but also strictly conformally Kähler w.r.t. both orientations [7]. The rod structure has one turning point and two semi-infinite rods [15, §4.2]. However, since the manifold is topologically trivial, there are no 2-cycles; thus the parameter $n$ does not correspond to a charge integral like (1.3). For a different construction involving 3-cycles, see [13].
Example 4.3 (Taub-bolt).
The Taub-bolt instanton is obtained from (4.5) when $m=\frac{5}{4}n$. It is ALF, and the topology is $\mathbb{CP}^{2}\setminus\{{\rm pt.}\}$. We have $\Psi_{2}^{-}=n/[4(r-n)^{3}]$, $\Psi_{2}^{+}=9n/[4(r+n)^{3}]$, so both $W^{\pm}_{abcd}$ remain type D. The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.7]. As in Kerr we have two turning points, so $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1] and there is one independent 2-cycle, which we denote by $S_{2n}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=2n$. We then find:
$\int_{S_{2n}}\hat{\omega}^{\pm}=c^{\pm}\,n^{2/3}$ (4.8)
where $c^{-}=-3\cdot 4^{1/3}\pi$, $c^{+}=3^{1/3}\cdot 4^{1/3}\pi$.
Examples 4.1, 4.2 and 4.3 are AF/ALF. Our construction also applies to the ALE case:
Example 4.4 (Eguchi-Hanson).
The metric for the Eguchi-Hanson instanton is locally given by
$g=f(r)\frac{r^{2}}{4}({\rm d}\psi+\cos\theta{\rm d}\phi)^{2}+\frac{{\rm d}r^{2}}{f(r)}+\frac{r^{2}}{4}({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}),$ (4.9)
where $f(r)=1-(a/r)^{4}$, and $a$ is a real constant. The topology is $T^{*}\mathbb{CP}^{1}$, and the asymptotics is ALE. The metric is Ricci-flat, the self-dual Weyl tensor is $W^{+}_{abcd}=0$, and the anti-self-dual Weyl tensor $W^{-}_{abcd}$ is type D  [^4] As noted by Gibbons and Pope [27], the Eguchi-Hanson metric (4.9) can be obtained from (4.5) by setting $m=n+a^{4}/(128n^{3})$, $r=m+\tilde{r}^{2}/(8n)$, and taking the limit $n\to\infty$ (and then redefining $\tilde{r}\to r$), see [27, Eqs. (4.15)-(4.16)]. This procedure leads to (4.10) (i.e., $W^{-}_{abcd}\neq 0$ and $W^{+}_{abcd}=0$, which is the opposite to self-dual Taub-NUT).:
$\Psi_{2}^{-}=\frac{a^{4}}{4r^{6}},\qquad\Psi_{2}^{+}=0.$ (4.10)
One can check that the following 2-forms
$\omega^{\pm}=\frac{r}{2}({\rm d}\psi+\cos\theta{\rm d}\phi)\wedge{\rm d}r\pm\frac{r^{2}}{4}\sin\theta{\rm d}\theta\wedge{\rm d}\phi$
give integrable almost-complex structures $J^{\pm}$ with opposite orientation, and that ${\rm d}\omega^{+}=0$, so $J^{+}$ is Kähler. However ${\rm d}\omega^{-}\neq 0$, and this gives the conformally Kähler form $\hat{\omega}^{-}=(\Psi_{2}^{-})^{2/3}\omega^{-}$ needed to define the charge (1.3). The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.5]. As in previous cases, two turning points give $H_{2}(M)\cong\mathbb{Z}$ [36, Thm. 1] so there is one independent 2-cycle, which we denote by $S_{a}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=a$. We have
$\int_{S_{a}}\hat{\omega}^{-}=-\pi\left(\tfrac{a}{4}\right)^{2/3}.$ (4.11)
So far the examples are Ricci-flat, but we can also include a non-trivial cosmological constant $\lambda\neq 0$. For $\lambda>0$ the instanton is necessarily compact, and Einstein-Hermitian (non-Kähler) compact 4-manifolds with $\lambda>0$ have been classified [33]: these are the $\mathbb{CP}^{2}$ instanton [26] (with orientation opposite to the Kähler one), the Page metric on $\mathbb{CP}^{2}\#\overline{\mathbb{CP}}^{2}$ [37], or the Chen-LeBrun-Weber metric on $\mathbb{CP}^{2}\#2\overline{\mathbb{CP}}^{2}$ [14]. Only the first two metrics are explicitly known.
Example 4.5 ($\mathbb{CP}^{2}$).
The $\mathbb{CP}^{2}$ instanton [26] is
$g=\frac{r^{2}}{4f(r)^{2}}({\rm d}\psi+\cos\theta{\rm d}\phi)^{2}+\frac{{\rm d}{r}^{2}}{f(r)^{2}}+\frac{r^{2}}{4f(r)}({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}),$ (4.12)
where $f(r)=1+\frac{\lambda}{6}r^{2}$. The metric is Einstein, with cosmological constant $\lambda$. One can check that the following 2-forms
$\omega^{\pm}=\frac{r}{2f^{2}}({\rm d}\psi+\cos\theta{\rm d}\phi)\wedge{\rm d}{r}\pm\frac{r^{2}}{4f}\sin\theta{\rm d}\theta\wedge{\rm d}\phi$
give integrable almost-complex structures with opposite orientation. Furthermore, ${\rm d}\omega^{+}=0$, ${\rm d}\omega^{-}\neq 0$, and ${\rm d}\hat{\omega}^{-}=0$, where $\hat{\omega}^{-}=\frac{4f^{2}}{r^{4}}\omega^{-}$. Thus, (4.12) is Kähler on the self-dual side, and strictly conformally Kähler on the anti-self-dual side. We also have
$\Psi_{2}^{-}=0,\qquad\Psi_{2}^{+}=\frac{\lambda}{3}.$ (4.13)
The rod structure has three turning points [15], at $r=0$ and at $(r,\theta)=(\infty,0),(\infty,\pi)$. Thus the Euler characteristic is $\chi=3$, and since the manifold is compact, this gives $b_{2}(M)=1$, so there is only one 2-cycle, corresponding to the set $r=\infty$, $0<\theta<\pi$, which is a bolt of $\partial_{\psi}$ that we denote $S_{\infty}$. Since ${\rm d}\hat{\omega}^{-}=0$, a charge like (1.3) can be associated to $\hat{\omega}^{-}$; however, $\Psi_{2}^{-}=0$ implies that this would not be related to the gravitational field. For $\hat{\omega}^{+}=(\Psi_{2}^{+})^{2/3}\omega^{+}$, the charge is
$\int_{S_{\infty}}\hat{\omega}^{+}=(2\cdot 3^{1/3}\pi)\,\lambda^{1/3}.$ (4.14)
Example 4.6 (Page).
The Page metric [37, 24, 38] is given by
$g=\frac{3(1+\nu^{2})}{\lambda}$ $\left[\frac{f(r)\sin^{2}r}{4(3+\nu^{2})^{2}}({\rm d}\psi+\cos\theta{\rm d}\phi)^{2}+\frac{{\rm d}{r}^{2}}{f(r)}\right.$ (4.15)
$\left.+\frac{(1-\nu^{2}\cos^{2}r)}{(3+6\nu^{2}-\nu^{4})}({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2})\right]$
where $\nu$ is the positive root of $\nu^{4}+4\nu^{3}-6\nu^{2}+12\nu-3=0$ ($\nu\approx 0.2817$), and
$f(r)=\frac{3-\nu^{2}-\nu^{2}(1+\nu^{2})\cos^{2}r}{1-\nu^{2}\cos^{2}r}.$
The space is Einstein, with cosmological constant $\lambda$. Both Weyl tensors $W^{\pm}_{abcd}$ are type D, and a calculation shows that
$\Psi_{2}^{\pm}=\frac{\lambda}{3}\frac{(1-\nu^{2})^{2}}{(1+\nu^{2})}\frac{1}{(1\pm\nu\cos r)^{3}}.$ (4.16)
One can check that the 2-forms
$\omega^{\pm}=\frac{3(1+\nu^{2})}{\lambda(3+\nu^{2})}\left[\frac{\sin r}{2}({\rm d}\psi+\cos\theta{\rm d}\phi)\wedge{\rm d}r\pm\frac{(1-\nu^{2}\cos^{2}r)}{4\nu}\sin\theta{\rm d}\theta\wedge{\rm d}\phi\right]$
give oppositely oriented integrable almost-complex structures. The topology is $\mathbb{CP}^{2}\#\overline{\mathbb{CP}}^{2}$, and it corresponds to blowing up the nut of $\mathbb{CP}^{2}$ at $r=0$. The rod structure has four turning points [15] (at $(r,\theta)=(0,0),(0,\pi),(\pi,0),(\pi,\pi)$), so the Euler characteristic is 4. This gives $b_{2}(M)=2$, hence there are two independent 2-cycles. These are represented by the two bolts of $\partial_{\psi}$ (cf. [38]), which we denote by $S_{0},S_{\pi}$ and correspond to the sets $r=0,\pi$, $0<\theta<\pi$. The charges (1.3) are
$\int_{S_{r_{o}}}\hat{\omega}^{\pm}=c^{\pm}_{r_{o}}\,\lambda^{-1/3}$ (4.17)
where $r_{o}=0,\pi$, and
$c^{\pm}_{r_{o}}=\pm 3^{1/3}\frac{(1+\nu^{2})^{1/3}(1-\nu^{2})^{4/3}}{\nu(3+\nu^{2})}\frac{(1\mp\nu\cos r_{o})}{(1\pm\nu\cos r_{o})}\,\pi.$
4.2. The Chen-Teo instanton
The Chen-Teo family of ALF Ricci-flat metrics is [17, Eq. (2.1)]
$g=\frac{(F{\rm d}\tau+G{\rm d}\phi)^{2}}{(x-y)HF}+\frac{kH}{(x-y)^{3}}\left(\frac{{\rm d}{x}^{2}}{X}-\frac{{\rm d}{y}^{2}}{Y}-\frac{XY}{kF}{\rm d}\phi^{2}\right)$ (4.18)
where $X=P(x)$, $Y=P(y)$, $F=y^{2}X-x^{2}Y$, and
$H={}$ $(\nu x+y)[(\nu x-y)(a_{1}-a_{3}xy)-2(1-\nu)(a_{0}-a_{4}x^{2}y^{2})],$
$G={}$ $(\nu^{2}a_{0}+2\nu a_{3}y^{3}+a_{4}(2\nu-1)y^{4})X+((1-2\nu)a_{0}-2\nu a_{1}x-\nu^{2}a_{4}x^{4})Y,$
$P(u)={}$ $a_{0}+a_{1}u+a_{2}u^{2}+a_{3}u^{3}+a_{4}u^{4}.$ (4.19)
Here, $a_{0},...,a_{4},k,\nu$ are real constants, with $\nu\in[-1,1]$. It was shown in [2] that (4.18) is one-sided type D, and thus conformally Kähler only on one side (which we take to be the self-dual side). The Kähler form is $\hat{\omega}=\Psi_{2}^{2/3}\omega$ (recall (2.16)), and to find an explicit expression we follow [8]. After replacing eq. (5.3) into eq. (2.10) in this reference, we get
$\omega=\frac{\sqrt{k}}{(x-y)^{2}}$$\left\{(x{\rm d}{y}-y{\rm d}{x})\wedge{\rm d}\tau+\left(\frac{Gx}{F}+\frac{HyX}{F(x-y)}\right){\rm d}{y}\wedge{\rm d}\phi\right.$ (4.20)
$\left.-\left(\frac{Gy}{F}+\frac{HxY}{F(x-y)}\right){\rm d}{x}\wedge{\rm d}\phi\right\}$
The scalar $\Psi_{2}$ can be computed using [8, Eq. (2.28)]; we find
$\Psi_{2}=-\frac{(\nu+1)}{2k}\left(\frac{x-y}{\nu x+y}\right)^{3},$ (4.21)
which coincides with the expression deduced from [2, Eqs. (3.28) and (3.39)]. The closed Kähler form is then $\hat{\omega}=\Psi_{2}^{2/3}\omega$.
The vector fields $\partial_{\tau},\partial_{\phi}$ are Killing, and the solution is toric. The rod structure has three turning points, two finite rods and two semi-infinite rods [17, 2, 31]. These can be conveniently described in terms of the four roots, $r_{1},r_{2},r_{3},r_{4}$, of the polynomial $P$ in (4.19). Following Chen and Teo [17, §3.3], we have
  • $\mathcal{R}_{1}$: $x=r_{2}$, $r_{1}<y<r_{2}$, direction $\ell_{1}=(\alpha_{1},\beta_{1})$,
  • $\mathcal{R}_{2}$: $r_{2}<x<r_{3}$, $y=r_{1}$, direction $\ell_{2}=(\alpha_{2},\beta_{2})$,
  • $\mathcal{R}_{3}$: $x=r_{3}$, $r_{1}<y<r_{2}$, direction $\ell_{3}=(\alpha_{3},\beta_{3})$,
  • $\mathcal{R}_{4}$: $r_{2}<x<r_{3}$, $y=r_{2}$, direction $\ell_{4}=(\alpha_{4},\beta_{4})$,
where the notation $\ell_{i}=(\alpha_{i},\beta_{i})$ means $\ell_{i}=\alpha_{i}\partial_{\tau}+\beta_{i}\partial_{\phi}$, and the expressions for $\alpha_{i},\beta_{i}$ can be found in [17, Eqs. (3.14)-(3.15)]. It is convenient to adapt the Killing coordinates to the rod vectors $\ell_{1},\ell_{2}$ (cf. [2]). This is done by defining $(\tilde{\tau},\tilde{\phi})$ via $\tau=\alpha_{2}\tilde{\tau}+\alpha_{1}\tilde{\phi}$, $\phi=\beta_{2}\tilde{\tau}+\beta_{1}\tilde{\phi}$, so that $\partial_{\tilde{\phi}}=\ell_{1}$ and $\partial_{\tilde{\tau}}=\ell_{2}$. The closed Kähler form $\hat{\omega}=\Psi_{2}^{2/3}\omega$ is then
$\hat{\omega}=\left(\tfrac{\nu+1}{2k}\right)^{2/3}\frac{\sqrt{k}}{(\nu x+y)^{2}}$$\left\{\frac{1}{\kappa_{2}}\left[\left(K_{2}[1]+\frac{G}{F}\right)x+\frac{HyX}{F(x-y)}\right]{\rm d}{y}\wedge{\rm d}\tilde{\tau}\right.$ (4.22)
$-\frac{1}{\kappa_{2}}\left[\left(K_{2}[1]+\frac{G}{F}\right)y+\frac{HxY}{F(x-y)}\right]{\rm d}{x}\wedge{\rm d}\tilde{\tau}$
$+\frac{1}{\kappa_{1}}\left[\left(K_{1}[1]+\frac{G}{F}\right)x+\frac{HyX}{F(x-y)}\right]{\rm d}{y}\wedge{\rm d}\tilde{\phi}$
$\left.-\frac{1}{\kappa_{1}}\left[\left(K_{1}[1]+\frac{G}{F}\right)y+\frac{HxY}{F(x-y)}\right]{\rm d}{x}\wedge{\rm d}\tilde{\phi}\right\},$
where we used the notation $\alpha_{i}=K_{i}[1]/\kappa_{i}$ and $\beta_{i}=1/\kappa_{i}$ following [17, Eqs. (3.14)-(3.15)].
The family of ALF metrics (4.18) has, in general, conical singularities. Chen and Teo showed in [17, §4.3] that, by imposing certain constraints on the parameters, one can obtain a smooth, AF, regular 4-manifold. This is the Chen-Teo instanton [16], which corresponds to setting the total NUT charge to zero and requiring the rod vectors to have closed orbits with $2\pi$ period and to satisfy $\ell_{1}=\ell_{4}=\pm\ell_{2}\pm\ell_{3}$, cf. [17, Eq. (4.13)] and [2]. The solution to these constraints is given in [17, Eq. (4.16)], from where one deduces that
$\nu\equiv-2\xi^{2},$
$r_{1}=-4\xi^{3}(1-\xi),\quad r_{2}=-\xi(1-2\xi+2\xi^{2}),\quad r_{3}=1-2\xi,\quad r_{4}=\infty,$
$a_{4}=0,\quad a_{3}=1,\quad a_{2}=-1+3\xi-2\xi^{2}+6\xi^{3}-4\xi^{4},$
$a_{1}=-\xi+4\xi^{2}-10\xi^{3}+20\xi^{4}-20\xi^{5}+16\xi^{6}-8\xi^{7},$
$a_{0}=4(1-\xi)(2\xi-1)(1-2\xi+2\xi^{2})\xi^{4}.$ (4.23)
The resulting family of metrics has 2 parameters, $(\xi,k)$. The parametrization (4.23) gives $\alpha_{1}=\alpha_{4}$, $\beta_{1}=\beta_{4}$, $\alpha_{3}=-(\alpha_{1}+\alpha_{2})$ $\beta_{3}=-(\beta_{1}+\beta_{2})$. The rod vectors are then $\ell_{1}=\ell_{4}=\partial_{\tilde{\phi}}$, $\ell_{2}=\partial_{\tilde{\tau}}$, $\ell_{3}=-(\partial_{\tilde{\phi}}+\partial_{\tilde{\tau}})$.
The topology is $\mathbb{CP}^{2}\setminus S^{1}$. Since there are three turning points, we have $H_{2}(M)\cong\mathbb{Z}^{2}$ [36, Thm. 1], so there are two independent 2-cycles, which we can take to be the finite rods $\mathcal{R}_{2},\mathcal{R}_{3}$. Using the expression (4.22), it is straightforward to integrate $\hat{\omega}$ over $\mathcal{R}_{2},\mathcal{R}_{3}$. Since the rods correspond to roots of $X,Y$, the terms with $H$ in (4.22) vanish on the rods, and also (from (4.19)) $G/F$ becomes constant, so we find (using $a_{4}=0$, $a_{3}=1$)
$\hat{\omega}\big{|}_{\mathcal{R}_{2}}={}$ $-\left(\tfrac{\nu+1}{2k}\right)^{2/3}\tfrac{\sqrt{k}}{\kappa_{1}r_{1}}\left(K_{1}[1]r_{1}^{2}+\nu^{2}a_{0}+2\nu r_{1}^{3}\right)\frac{{\rm d}{x}\wedge{\rm d}\tilde{\phi}}{(\nu x+r_{1})^{2}},$
$\hat{\omega}\big{|}_{\mathcal{R}_{3}}={}$ $\left(\tfrac{\nu+1}{2k}\right)^{2/3}\tfrac{\sqrt{k}}{r_{3}}\left[\left(\tfrac{K_{2}[1]}{\kappa_{2}}-\tfrac{K_{1}[1]}{\kappa_{1}}\right)r_{3}^{2}\right.$
$\left.-(\tfrac{1}{\kappa_{2}}-\tfrac{1}{\kappa_{1}})[(1-2\nu)a_{0}-2\nu a_{1}r_{3}]\right]\frac{{\rm d}{y}\wedge{\rm d}\tilde{\tau}}{(\nu r_{3}+y)^{2}}$
where for $\hat{\omega}\big{|}_{\mathcal{R}_{3}}$ we used the fact that $\tilde{\tau}+\tilde{\phi}$ is constant on $\mathcal{R}_{3}$. Integrating and replacing (4.23), we find the charges
$\int_{\mathcal{R}_{2}}\hat{\omega}={}$ $\frac{4\pi}{2^{2/3}}\left(\frac{k}{1-2\xi^{2}}\right)^{1/3},$ (4.24)
$\int_{\mathcal{R}_{3}}\hat{\omega}={}$ $-\frac{16\pi}{2^{2/3}}\left(\frac{k}{1-2\xi^{2}}\right)^{1/3}\xi^{2}.$ (4.25)
Remark 4.7.
This result implies that both of the two parameters of the Chen-Teo instanton can be obtained as charge integrals of the Kähler form. As mentioned in the introduction, this is in contrast to the Kerr case, where the Kähler construction gives only one of the two parameters, namely the mass, see (4.4). We can understand this difference from the fact that there are two bolts in Chen-Teo, whereas Kerr has only one. In a sense, the Chen-Teo instanton is a “double” object, as suggested by the interpretation in [17, §5] in terms of two touching Kerr-NUTs.
5. Linearized Kähler structures
A compact Einstein metric with the property $\det(W^{+})>0$ is necessarily conformally Kähler, a remarkable result shown by Wu in [42]. LeBrun’s alternative proof [34] can be adapted to manifolds with a different asymptotic structure as in [12, Proposition 3]. This in turn was used in [12] to show that Ricci-flat metrics sufficiently close to a Ricci-flat, ALF, toric, Hermitian instanton must also be Hermitian.
In this section we shall use ideas from [34, 12] to deduce some identities that allow us to interpret our basic result (3.20) as the linearized version of a complex structure for the perturbed instanton.
5.1. A divergence identity
As in [34, 12], starting from a Riemannian manifold $(M,g_{ab})$, it turns out to be convenient to work with a conformally transformed metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$. Below, our convention is that indices of a hatted quantity are moved with $\hat{g}_{ab},\hat{g}^{ab}$ (and spinor versions), and indices of an unhatted quantity are moved with $g_{ab}$, etc.
We recall that the eigenvalue problem for the Weyl spinor $\Psi_{AB}{}^{CD}$ and its conformal transformation $\hat{\Psi}_{AB}{}^{CD}$ was summarized in sections 2.2 and 2.3. Let $\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\lambda}_{3}$ be the eigenvalues of $\hat{\Psi}_{AB}{}^{CD}$. Assume $\hat{\lambda}_{3}\neq 0$, and let $\hat{\varphi}_{AB}$ be an eigenspinor with unit norm corresponding to $\hat{\lambda}_{3}$,
$\hat{\Psi}_{AB}{}^{CD}\hat{\varphi}_{CD}=\hat{\lambda}_{3}\,\hat{\varphi}_{AB},\qquad|\hat{\varphi}|_{\hat{g}}^{2}=1,$ (5.1)
where, here and in what follows, $|\cdot|_{\hat{g}}^{2}$ denotes the norm with respect to $\hat{g}_{ab}$, $\hat{I}_{ABCD}$ and their inverses, so for example $|\hat{j}|_{\hat{g}}^{2}=\hat{g}_{ab}\hat{j}^{a}\hat{j}^{b}$ for any $\hat{j}^{a}$, and
$|\hat{\varphi}|_{\hat{g}}^{2}=\hat{I}^{ABCD}\hat{\varphi}_{AB}\hat{\varphi}_{CD},\qquad|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}=\hat{g}^{ab}\hat{I}^{CDEF}\hat{\nabla}_{a}\hat{\varphi}_{CD}\hat{\nabla}_{b}\hat{\varphi}_{EF}.$ (5.2)
The following result is based on the constructions of Wu [42] and LeBrun [34]. It is a slight generalization in that the main formula (5.4) below is an identity valid for generic Riemannian 4-manifolds; in particular, it does not assume the Einstein condition, compactness, or a specific sign of the determinant of the Weyl tensor.
Lemma 5.1.
Let $(M,\hat{g}_{ab})$ be a Riemannian 4-manifold, and let $\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\lambda}_{3}$ be the eigenvalues of $\hat{\Psi}_{AB}{}^{CD}$. Assume $\hat{\lambda}_{3}\neq 0$ in an open neighbourhood, and let $\hat{\varphi}_{AB}$ be a unit eigenspinor corresponding to $\hat{\lambda}_{3}$, eq. (5.1). Define the vector fields
$\hat{j}^{AA^{\prime}}=\hat{\nabla}^{A^{\prime}B}\hat{\varphi}^{A}{}_{B},\qquad\hat{V}^{AA^{\prime}}=\hat{\varphi}^{AC}\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC}.$ (5.3)
Then the following identity holds
^$\hat{\nabla}_{a}\hat{V}^{a}={}$ $\frac{1}{6}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+|\hat{j}|_{\hat{g}}^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2})\right]$ (5.4)
$+\frac{2}{3}\hat{\varphi}^{AB}\hat{\varphi}^{CD}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{A^{\prime}E}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})$
where $\hat{A}_{a}$ and $\hat{B}_{a}$ are the 1-forms (2.5a), (2.5b) with respect to $\hat{g}_{ab}$.
Remark 5.2.
The spinor objects in the first line of (5.4) can be written in tensor terms by introducing the self-dual 2-form ^$\hat{\mathcal{F}}_{ab}=\sqrt{2}\,\hat{\varphi}_{AB}\hat{\epsilon}_{A^{\prime}B^{\prime}}$,
$\hat{V}^{a}=\frac{1}{2}\hat{\mathcal{F}}^{ac}\hat{\nabla}^{b}\hat{\mathcal{F}}_{bc},\qquad\hat{j}_{a}=\frac{1}{\sqrt{2}}\hat{\nabla}^{b}\hat{\mathcal{F}}_{ba},\qquad|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}=\frac{1}{4}|\hat{\nabla}\hat{\mathcal{F}}|_{\hat{g}}^{2}.$ (5.5)
Proof.
This is based on a series of calculations, which for convenience we summarized in the form of two propositions B.1 and B.2 in Appendix B. Referring to the proof of these propositions: eq. (B.4) gives an expression for $\hat{\lambda}_{3}-\frac{\hat{R}}{6}$; replacing this in (B.2), we immediately get (5.4). ∎
Recall that the identity (5.4) is formulated in terms of the metric $\hat{g}_{ab}=\Omega^{2}g_{ab}$. To understand the last term in (5.4), we write it in terms of $g_{ab}$,
^$\hat{\nabla}^{AA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})=\hat{\lambda}_{3}^{-3}\nabla^{AA^{\prime}}\Psi_{ABCD},$ (5.6)
which implies that it vanishes if $(M,g_{ab})$ is Einstein. In the spirit of Wu [42] and LeBrun [34], it follows in this case, that in view of (3.8) and (5.6), the second line in (5.4) vanishes. Hence, integrating (5.4), taking this vanishing into account, and using the divergence theorem, we have
$\oint_{\partial M}\hat{V}\lrcorner\,\hat{\boldsymbol{\epsilon}}=\frac{1}{6}\int_{M}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}\,\hat{\boldsymbol{\epsilon}}+\hat{P},$ (5.7)
where $\hat{\boldsymbol{\epsilon}}$ is the volume form of $\hat{g}_{ab}$ and
$\hat{P}=\int_{M}\left(|\hat{j}|_{\hat{g}}^{2}+\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2})\right]\right)\hat{\boldsymbol{\epsilon}}.$ (5.8)
Thus, if the boundary term vanishes and $\hat{P}$ is positive, then $(M,\hat{g}_{ab})$ is necessarily Kähler, since the equation |^$|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}=0$ implies ^$\hat{\nabla}_{a}\hat{\varphi}_{BC}=0$, so $J^{a}{}_{b}=\hat{\mathcal{F}}_{bc}\hat{g}^{ca}$ is a parallel almost-complex structure. The Einstein (or Ricci-flat) manifold $(M,g_{ab})$ is then conformally Kähler.
As mentioned, for compact Einstein manifolds with $\det(W^{+})>0$ the above result was shown by Wu [42]. Here, we followed LeBrun’s formulation [34] but dropping the compactness assumption (as in [12]) and the conditions $\nabla\cdot W^{+}=0$ and $\det(W^{+})>0$. The fact that $\det(W^{+})>0$ ensures $\hat{P}\geq 0$ follows from (2.10), since (2.10a) implies that one can choose $\hat{\lambda}_{3}>0$ and $\hat{\lambda}_{1}<0$, so, from (2.10c), ${\rm sign}\det(W^{+})=-{\rm sign}\,\hat{\lambda}_{2}$.
5.2. Perturbative expansions
If we do not assume that the boundary term in (5.7) vanishes or that $\hat{P}$ is positive, then we cannot conclude that $(M,\hat{g}_{ab})$ is Kähler. However, our goal now is to study infinitesimal Einstein deformations of (5.7) around a Hermitian instanton. We recall that our notation and treatment for perturbations was summarized in section 3.2. A perturbative expansion of the eigenvalue problem of the map $\hat{W}^{+}_{ab}{}^{cd}$ is justified by the fact that the latter is real and symmetric, see [30].
Theorem 5.3.
Let $g_{ab}(s)$ be a smooth, one-parameter family of Einstein metrics in $M$, and consider the conformally related metrics $\hat{g}_{ab}(s)$ as in section 5.1. Suppose that $g_{ab}=g_{ab}(0)$ is Hermitian (so $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, with Kähler form ^$\hat{\mathcal{F}}_{ab}$). Then the perturbative expansion of the identities (5.4) and (5.7) around $s=0$ gives the following formulas
${\rm d}\,\delta\hat{\mathcal{F}}={}$ $0,$ (5.9)
$\frac{1}{4}\oint_{\partial M}\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\,(\delta^{2}\hat{\mathcal{F}})={}$ $\int_{M}\left[\frac{1}{24}\big{|}\delta(\hat{\nabla}\hat{\mathcal{F}})\big{|}_{\hat{g}}^{2}+\big{|}\delta\hat{j}\big{|}_{\hat{g}}^{2}\right.$
$\left.+\frac{1}{3\hat{\lambda}_{3}}\left((\delta\hat{\lambda}_{1}-\delta\hat{\lambda}_{2})^{2}+\hat{\lambda}_{3}(|\delta{\hat{A}}|_{\hat{g}}^{2}+|\delta\hat{B}|_{\hat{g}}^{2})\right)\right]\hat{\boldsymbol{\epsilon}}$ (5.10)
Furthermore, if the left hand side of (5.10) vanishes, then
$\left.\frac{{\rm d}}{{\rm d}{s}}\left[\hat{\nabla}_{a}^{(s)}\hat{\mathcal{F}}_{bc}(s)\right]\right|_{s=0}=0.$ (5.11)
In particular, (5.11) holds for perturbations of compact instantons with $\lambda>0$ and for ALF vacuum perturbations of ALF Ricci-flat instantons.
Proof.
We first analyze the perturbative expansion of the right hand side of (5.4). Since $\hat{g}_{ab}=\hat{g}_{ab}(0)$ is Kähler, the background self-dual Weyl tensor $\hat{W}^{+}_{abcd}$ has two repeated eigenvalues: $\hat{\lambda}_{1}=\hat{\lambda}_{2}=-\hat{\lambda}_{3}/2$, where $\hat{\lambda}_{3}>0$. The parallel background Kähler form is ^$\hat{\mathcal{F}}_{ab}$. Recalling (5.5), we then have:
$\hat{g}_{ab}(s)={}$ $\hat{g}_{ab}+\delta\hat{g}_{ab}\,s+O(s^{2}),$
$\hat{j}_{a}(s)={}$ $\delta\hat{j}_{a}\,s+O(s^{2}),$
^(s)^$\hat{\nabla}^{(s)}\hat{\mathcal{F}}(s)={}$ $\delta(\hat{\nabla}\hat{\mathcal{F}})\,s+O(s^{2}),$
$\hat{\lambda}_{1}(s)={}$ $-\hat{\lambda}_{3}/2+\delta\hat{\lambda}_{1}\,s+O(s^{2}),$
$\hat{\lambda}_{2}(s)={}$ $-\hat{\lambda}_{3}/2+\delta\hat{\lambda}_{2}\,s+O(s^{2}),$
$\hat{A}_{a}(s)={}$ $\delta\hat{A}_{a}\,s+O(s^{2}),$
$\hat{B}_{a}(s)={}$ $\delta\hat{B}_{a}\,s+O(s^{2}).$
Replacing in (5.4), we see that the first non-trivial term is $O(s^{2})$, and is equal to the integrand on the right hand side of (5.10).
For a perturbative expansion of the left hand side of (5.4) (the divergence term), it is convenient to multiply by the volume form $\hat{\boldsymbol{\epsilon}}$, so that we can use the identity (^$(\hat{\nabla}_{a}\hat{V}^{a})\,\hat{\boldsymbol{\epsilon}}=-{\rm d}(\hat{*}\hat{V})$, where on the right side, $\hat{V}$ is interpreted as a 1-form. Using (5.5), one can also check that ^$\hat{*}\hat{V}=-\frac{1}{2}\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}$. Thus:
(^$(\hat{\nabla}_{a}\hat{V}^{a})\,\hat{\boldsymbol{\epsilon}}=\tfrac{1}{2}{\rm d}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}).$
The perturbative expansion of ^^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}}$ is:
(^^$(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})(s)=\delta(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,s+\frac{1}{2!}\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})\,s^{2}+O(s^{3}),$
where
$\delta(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})={}$^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta\hat{\mathcal{F}}),$
$\delta^{2}(\hat{\mathcal{F}}\wedge\hat{*}{\rm d}\hat{\mathcal{F}})={}$^$\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta^{2}\hat{\mathcal{F}})+2\delta\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta\hat{\mathcal{F}})+2\hat{\mathcal{F}}\wedge(\delta\hat{*}){\rm d}(\delta\hat{\mathcal{F}}).$
Since there is no $O(s)$ term on the right side of (5.7), equation (5.9) follows. Equality of the $O(s^{2})$ terms gives (5.10).
To prove (5.11), note that every term on the right hand side of (5.10) is positive definite, so if the boundary term on the left vanishes, then every term on the right side vanishes as well. In particular $\delta(\hat{\nabla}\hat{\mathcal{F}})=0$, which is eq. (5.11).
Finally, we analyze the boundary term. In the compact case with $\lambda>0$, there is no boundary, so the left hand side of (5.10) vanishes. For the ALF case, we write the 3-form in (5.10) in indices,
[^^$[\hat{\mathcal{F}}\wedge\hat{*}{\rm d}(\delta^{2}\hat{\mathcal{F}})]_{abc}=\frac{3}{2}\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}.$ (5.12)
We need to integrate this form over a 3-surface $\Sigma_{r}=\{r=const.\}$ (see [10, 3] for the definition of $r$), and then take the limit $r\to\infty$. In the following, all computations will be done with respect to the metric $g_{ab}$. In particular, for any tensor field $t_{a_{1}...a_{n}}$, the norm $|t|_{g}$ is defined by
$|t|_{g}:=(g^{a_{1}b_{1}}...g^{a_{n}b_{n}}t_{a_{1}...a_{n}}t_{b_{1}...b_{n}})^{1/2}.$ (5.13)
Let $\alpha_{abc}$ denote the 3-form (5.12) restricted to $\Sigma_{r}$, then $\alpha_{abc}=f\mu_{abc}$ for some function $f$, where $\mu_{abc}$ is the volume 3-form on $\Sigma_{r}$ induced from the volume 4-form of $g_{ab}$. Using (5.13) we have $|\mu|_{g}=\sqrt{3!}$, hence $|f|=\frac{1}{\sqrt{3!}}|\alpha|_{g}$. To estimate $|\alpha|_{g}$, we first use that ^$\hat{\mathcal{F}}_{ab}=\Omega^{2}\mathcal{F}_{ab}$, $\hat{\epsilon}_{abcd}=\Omega^{4}\epsilon_{abcd}$, $\hat{g}^{ab}=\Omega^{-2}g^{ab}$, and that $|\mathcal{F}|_{g}=\sqrt{2}$, $|\epsilon|_{g}=\sqrt{4!}$, $|g|_{g}=2$, then
|^$|\hat{\mathcal{F}}|_{g}=\sqrt{2}\,\Omega^{2},\qquad|\hat{\epsilon}|_{g}=\sqrt{4!}\,\Omega^{4},\qquad|\hat{g}^{-1}|_{g}=2\,\Omega^{-2}.$ (5.14)
Here, $\Omega=\lambda_{3}^{1/3}$. From [10], the ALF condition implies that
$\Omega=\frac{1}{r}+\text{lower order}.$ (5.15)
Combining (5.14) and (5.15), we get ^$\hat{\mathcal{F}}_{ab}=O_{g}(r^{-2})$, $\hat{\epsilon}_{abcd}=O_{g}(r^{-4})$, and $\hat{g}^{ab}=O_{g}(r^{2})$, where we recall that the notation $t_{a_{1}...a_{n}}=O_{g}(r^{\alpha})$ means that there is a constant $C$ such that $|t|_{g}\leq Cr^{\alpha}$ for sufficiently large $r$, see [3, Def. 2.6]. Thus, the factor ^$\hat{\mathcal{F}}_{[ab}\hat{\epsilon}_{c]klm}\hat{g}^{kd}\hat{g}^{le}\hat{g}^{mf}$ in (5.12) is $O_{g}(1)$. To estimate the remaining factor $\partial_{d}(\delta^{2}\hat{\mathcal{F}})_{ef}$, we first use that $\delta^{2}\hat{\mathcal{F}}=\delta^{2}(\Omega^{2})\mathcal{F}+\Omega^{2}\delta^{2}\mathcal{F}+2\delta(\Omega^{2})\delta\mathcal{F}$, so every term on the right side is $O_{g}(r^{-2})$. In addition, due to the skew-symmetry of (5.12), $\partial_{d}$ can be replaced by the Levi-Civita connection $\nabla_{d}^{(s)}$of $g_{ab}(s)$, so we can use that by assumption each derivative has extra fall-off, hence $\nabla_{d}^{(s)}(\delta^{2}\hat{\mathcal{F}})_{ef}=O_{g}(r^{-3})$.
Summarizing, so far we have that the 3-form (5.12) on $\Sigma_{r}$ is $\alpha_{abc}\propto|\alpha|_{g}\mu_{abc}$ and $|\alpha|_{g}=O_{g}(r^{-3})$. From [10, Def. 1.1], $\mu_{abc}$ goes like $r^{2}$. Hence, the integrand in the left hand side of (5.10) is order $r^{-1}$, thus it vanishes in the limit $r\to\infty$. ∎
Remark 5.4.
To relate $\delta\hat{\mathcal{F}}_{ab}$ to the closed 2-form $\delta\hat{\omega}_{ab}$ of section 3.2, choose $Z_{ab}(s)$ to be an eigenform of $W^{+}_{ab}{}^{cd}(s)$ in (3.17). We get $\Sigma_{ab}(s)=-2\Psi_{2}(s)Z_{ab}(s)$, and
$\delta\hat{\omega}_{ab}={}$ $-\frac{\sqrt{2}}{3}\left[\Psi_{2}^{-1/3}\delta\Sigma_{ab}-\Psi_{2}^{2/3}\delta{Z}_{ab}\right]$
[…]
so the result (5.9) is actually the same as (3.20).
Eq. (5.11) can be written more explicitly in terms of perturbed tensor fields, by using the perturbative expansions given in section 3.2. Recalling formulas (3.14) and (3.15) (formulated in terms of $\hat{g}_{ab}$), and using ^$\hat{\nabla}_{a}\hat{\mathcal{F}}_{bc}=0$, we have
$\left.\frac{{\rm d}}{{\rm d}{s}}\left[\hat{\nabla}_{a}^{(s)}\hat{\mathcal{F}}_{bc}(s)\right]\right|_{s=0}=\hat{\nabla}_{a}\delta\hat{\mathcal{F}}_{bc}-\delta\hat{C}_{ab}{}^{d}\hat{\mathcal{F}}_{dc}-\delta\hat{C}_{ac}{}^{d}\hat{\mathcal{F}}_{bd},$ (5.17)
where
$\delta\hat{C}_{ab}{}^{d}=\frac{1}{2}\hat{g}^{de}(\hat{\nabla}_{a}\delta\hat{g}_{be}+\hat{\nabla}_{b}\delta\hat{g}_{ae}-\hat{\nabla}_{e}\delta\hat{g}_{ab}).$ (5.18)
Note that (5.17)=0italic-(5.17italic-)$=0$ gives again ${\rm d}\,\delta\hat{\mathcal{F}}=0$.
5.3. Proof of Theorem 1.1
Let $g_{ab}(s)$ be a curve in $\mathscr{E}(M)$, the space of ALF Ricci-flat metrics on $M$, with $g_{ab}(0)=g_{ab}$, which is by assumption Hermitian non-Kähler. The tensor field $J^{a}{}_{b}(s)=\hat{\omega}_{bc}(s)\hat{g}^{ca}(s)$ is an almost-complex structure, and Theorem 5.3 implies that
^$\hat{\nabla}_{a}^{(s)}J^{b}{}_{c}(s)=O(s^{2}),$ (5.19)
i.e. it is parallel to second order. It follows that the curve $g_{ab}(s)$ is conformally Kähler (in particular, Hermitian) to second order. Hence the first order perturbation $\delta{g}_{ab}$ solves the linearized Hermitian-Einstein condition. Since Hermitian, ALF instantons have been classified [10, 35] and the moduli space is smooth and finite-dimensional, and since the topology of $M$ is fixed, it follows that $\delta{g}_{ab}$ must be a perturbation with respect to the moduli parameters, up to gauge. Thus, any ALF Ricci-flat perturbation is tangent to the moduli space of Hermitian metrics, in other words, infinitesimal rigidity and integrability hold.
Acknowledgements
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Twistor Theory where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1. BA is supported by the ERC Consolidator/UKRI Frontier grant TwistorQFT EP/Z000157/1. The work of LA is partially supported by the National Natural Science Foundation of China, under Grant Number W2431012.
Appendix A Conformal-GHP connections
Consider the self-dual 2-form $Z_{ab}\equiv Z^{3}_{ab}$ given in (2.1). Its principal spinors are $o_{A},\iota_{A}$. The tensor field $J^{a}{}_{b}\equiv\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Note that $J^{a}{}_{b}$ is invariant under the “GHP” transformation
$o_{A}\to zo_{A},\qquad\iota_{A}\to z^{-1}\iota_{A},$ (A.1)
where $z\in U(1)$. The almost-complex structure is also invariant under a conformal transformation
$g_{ab}\to\Omega^{2}g_{ab},\qquad o_{A}\to\Omega^{1/2}o_{A},\qquad\iota_{A}\to\Omega^{1/2}\iota_{A},$ (A.2)
A spinor/tensor field $\varphi$ which transforms under (A.1)-(A.2) as $\varphi\to\Omega^{2}z^{p}\varphi$ is said to have conformal weight $w$ and GHP weight $p$. These ‘weighted’ fields are to be thought of as sections of vector bundles over $M$. A linear connection $\mathcal{C}_{AA^{\prime}}$ on these bundles was constructed in [6]. If $\phi_{AB...K}$ is a totally symmetric spinor field with conformal weight $w$ and GHP weight $p$, we have:
$\mathcal{C}_{AA^{\prime}}\phi_{BC...L}={}$ $\nabla_{AA^{\prime}}\phi_{BC...L}+wf_{AA^{\prime}}\phi_{BC...L}+pP_{AA^{\prime}}\phi_{BC...L}$ (A.3)
$-f_{BA^{\prime}}\phi_{AC...L}-f_{CA^{\prime}}\phi_{BA...L}-...-f_{LA^{\prime}}\phi_{BC...A},$
where the 1-forms $f_{a}$ and $P_{a}$ are
$f_{a}=-\frac{1}{2}J^{c}{}_{b}\nabla_{c}J^{b}{}_{a},\qquad P_{a}=\frac{i}{2}(C_{a}-J^{b}{}_{a}f_{b}),$ (A.4)
and $C_{a}$ was defined in (2.5c). Below we collect a few facts needed in the main text:
  1. (1)
    If $\phi_{AB...K}$ has $w=-1$, $p=0$, then
    $\mathcal{C}_{A^{\prime}}{}^{B}\phi_{AB...K}=\nabla_{A^{\prime}}{}^{B}\phi_{AB...K},$ (A.5)
    whereas for $w=0$, $p=0$:
    $\mathcal{C}_{A^{\prime}}{}^{B}\phi_{AB...K}=\nabla_{A^{\prime}}{}^{B}\phi_{AB...K}+f_{A^{\prime}}{}^{B}\phi_{AB...K}.$ (A.6)
  2. (2)
    $o_{A}$ and $\iota_{A}$ have weights $(w=\frac{1}{2},p=1)$ and $(w=\frac{1}{2},p=-1)$ respectively, so a calculation using (A.3) gives:
    $\mathcal{C}_{AA^{\prime}}o_{B}=\sigma^{0}_{A^{\prime}}\iota_{A}\iota_{B},\qquad\sigma^{0}_{A^{\prime}}=o^{A}o^{B}\nabla_{AA^{\prime}}o_{B},$ (A.7a)
    $\mathcal{C}_{AA^{\prime}}\iota_{B}=\sigma^{1}_{A^{\prime}}o_{A}o_{B},\qquad\sigma^{1}_{A^{\prime}}=\iota^{A}\iota^{B}\nabla_{AA^{\prime}}\iota_{B}.$ (A.7b)
    Note that $\sigma^{1}_{A^{\prime}}=(\sigma^{0}_{A^{\prime}})^{\dagger}$.
  3. (3)
    $(M,g_{ab})$ is Hermitian w.r.t. $J^{a}{}_{b}$ if and only if $\sigma^{0}_{A^{\prime}}=0$.
Appendix B Some details of calculations
In this appendix we collect some details of calculations needed for section 5 in the main text.
Proposition B.1.
Let $(M,\hat{g}_{ab})$ be a Riemannian manifold, and let $\hat{\varphi}_{AB}$ be an eigenspinor of $\hat{\Psi}_{AB}{}^{CD}$ with eigenvalue $\hat{\lambda}_{3}$, normalized as $|\hat{\varphi}|_{\hat{g}}^{2}=1$. Define the vector fields
$\hat{j}^{AA^{\prime}}=\hat{\nabla}^{A^{\prime}B}\hat{\varphi}^{A}{}_{B},\qquad\hat{V}^{AA^{\prime}}=\hat{\varphi}^{AC}\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC}.$ (B.1)
Then the following identity holds:
^$\hat{\nabla}_{a}\hat{V}^{a}=|\hat{j}|_{\hat{g}}^{2}-\frac{1}{2}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}+\hat{\lambda}_{3}-\frac{\hat{R}}{6}.$ (B.2)
Proof.
For any symmetric spinor $\hat{\varphi}_{AB}=\hat{\varphi}_{(AB)}$, defining $\hat{j}^{a},\hat{V}^{a}$ as in (B.1), we have:
^$\hat{\nabla}_{AA^{\prime}}\hat{V}^{AA^{\prime}}={}$ (^$(\hat{\nabla}_{AA^{\prime}}\hat{\varphi}^{AC})(\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC})+\hat{\varphi}^{AC}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{A^{\prime}B}\hat{\varphi}_{BC}$
[…]
where we are using the convention (2.7) for the action of the curvature operator ^$\widehat{\Box}_{AB}$. Now choose $\hat{\varphi}_{AB}$ to be a unit eigenspinor $\hat{\Psi}_{AB}{}^{CD}$ with eigenvalue $\hat{\lambda}_{3}$:
$\hat{\Psi}_{AB}{}^{CD}\hat{\varphi}_{CD}=\hat{\lambda}_{3}\,\hat{\varphi}_{AB},\qquad|\hat{\varphi}|_{\hat{g}}^{2}=1.$ (B.3)
Then $\hat{\varphi}^{AB}\hat{\nabla}_{c}\hat{\nabla}^{c}\hat{\varphi}_{AB}=-|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}$, so (B.2) follows. ∎
Proposition B.2.
Let $(M,\hat{g}_{ab})$ be a Riemannian manifold, and let $\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\lambda}_{3}$ be the eigenvalues of $\hat{\Psi}_{AB}{}^{CD}$. Assume $\hat{\lambda}_{3}\neq 0$ in an open neighbourhood, and let $\hat{\varphi}_{AB}$ be a unit eigenspinor corresponding to $\hat{\lambda}_{3}$. Then the following identity holds:
$\frac{2}{3}\hat{\varphi}^{AB}\hat{\varphi}^{CD}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})=\\ -\frac{2}{3}|\hat{\nabla}\hat{\varphi}|_{\hat{g}}^{2}-\frac{1}{3\hat{\lambda}_{3}}\left[(\hat{\lambda}_{1}-\hat{\lambda}_{2})^{2}-2(\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2})\right]+\hat{\lambda}_{3}-\frac{\hat{R}}{6},$ (B.4)
where $\hat{A}_{a},\hat{B}_{a}$ are given by (2.5a), (2.5b) w.r.t. the metric $\hat{g}_{ab}$.
Proof.
Using ^$\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}=\frac{1}{2}\delta_{A}^{E}\widehat{\Box}+\hat{\epsilon}^{EF}\widehat{\Box}_{AF}$, and the action of ^$\widehat{\Box}_{AF}$ given in (2.7), we have:
$2\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})={}$$\widehat{\Box}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})-\frac{\hat{R}}{2\hat{\lambda}_{3}}\hat{\Psi}_{ABCD}$
$+\frac{6}{\hat{\lambda}_{3}}\hat{\Psi}_{AEF(B}\hat{\Psi}^{EF}{}_{CD)}.$
The term in the second line can be rewritten using [39, Eq. (3.3.55)], after which we get
$2\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})={}$$\widehat{\Box}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})-\frac{\hat{R}}{2\hat{\lambda}_{3}}\hat{\Psi}_{ABCD}$
$+\frac{6}{\hat{\lambda}_{3}}\hat{\Psi}_{ABEF}\hat{\Psi}^{EF}{}_{CD}-\frac{2|\hat{\Psi}|_{\hat{g}}^{2}}{\hat{\lambda}_{3}}\hat{I}_{ABCD},$
where $|\hat{\Psi}|_{\hat{g}}^{2}=\hat{\Psi}_{ABCD}\hat{\Psi}^{ABCD}$ and we recall $\hat{I}_{ABCD}=\hat{\epsilon}_{A(C}\hat{\epsilon}_{|B|D)}$ (see (2.2)). Contracting with $\hat{\varphi}^{AB}\hat{\varphi}^{CD}$ and using (B.3):
$2\hat{\varphi}^{AB}\hat{\varphi}^{CD}\hat{\nabla}_{AA^{\prime}}\hat{\nabla}^{EA^{\prime}}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{BCDE})={}$ $\hat{\varphi}^{AB}\hat{\varphi}^{CD}\widehat{\Box}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})-\frac{\hat{R}}{2}$
$+6\hat{\lambda}_{3}-\frac{2|\hat{\Psi}|_{\hat{g}}^{2}}{\hat{\lambda}_{3}}$
For the terms in the second line of the last equation, note that
$6\hat{\lambda}_{3}-\frac{2|\hat{\Psi}|_{\hat{g}}^{2}}{\hat{\lambda}_{3}}={}$ $\frac{1}{\hat{\lambda}_{3}}(6\hat{\lambda}_{3}^{2}-2(\hat{\lambda}_{1}^{2}+\hat{\lambda}_{2}^{2}+\hat{\lambda}_{3}^{2}))$
[…]
where in the first line we used (2.10b), and in the last we used (2.10a). On the other hand, using (B.3) again we find:
$\hat{\varphi}^{AB}\hat{\varphi}^{CD}\widehat{\Box}(\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD})={}$ $\hat{\lambda}_{3}^{-1}\hat{\Psi}_{ABCD}\widehat{\Box}(\hat{\varphi}^{AB}\hat{\varphi}^{CD})$
[…]
Recalling (2.11b) and (2.5a), (2.5b), we have
$\hat{\Psi}_{ABCD}\hat{\nabla}^{e}\hat{\varphi}^{AB}\hat{\nabla}_{e}\hat{\varphi}^{CD}=\hat{\lambda}_{1}|\hat{A}|_{\hat{g}}^{2}+\hat{\lambda}_{2}|\hat{B}|_{\hat{g}}^{2}.$
Putting everything together, the result (B.4) follows. ∎